Determine the common stock for Bertinelli Corp. based on the following information: cash = $340,000; patents and copyrights = $630,000; accounts payable = $500,000; accounts receivable = $139,000; tangible net fixed assets = $4,200,000; inventory = $185,000; notes payable = $170,000; accumulated retained earnings = $1,295,000; long-term debt = $1,830,000.
Answer #1

TOTAL ASSETS = Cash + Accounts receivable + Inventory + tangible net fixed assets + patents and copyrights

TOTAL ASSETS = 340,000 + 139,000 + 185,000 + 4,200,000 + 630,000 = 5,494,000

Total Assets must always be equal to Total Liabilities and Equity.

TOTAL LIABILITIES & EQUITY = 5,494,000

TOTAL LIABILITIES & EQUITY = Accounts payable + Notes payable + Long term debt + Common stock + Accumulated retained earnings

5,494,000 = 500,000 + 170,000 + 1,830,000 + Common stock + 1,295,000

Common stock = 5,494,000 - (500,000 + 170,000 + 1,830,000 + 1,295,000) = 1,699,000

Answer : Common stock = 1,699,000 (Thumbs up please)